Walker Incorporated sold $500,000 of 10% bonds on January 1, 2020 for a price that yields a 12% interest rate. The bonds pay interest semi-annually on June 30 and December 31. The bonds are due December 31, 2024. Walker uses the effective interest method.

Instructions:

1. Determine the selling price of the bonds on January 1, 2020.

2. Prepare an amortization schedule using the effective interest method.

3. Prepare the journal entries for 2020.

4. Assume the company reacquired the bonds on July 1, 2023, at 104 and prepare journal entry to record the retirement of the bonds.

5. Assume Walker Incorporated used the straight-line method to amortize the premium or discount. Prepare the journal for June 30, 2020 to record interest expense.
Answer #1

Solution

Walker Incorporated

  1. Determination of the selling price of bonds on January 1, 2020:

Selling price of bonds = present value of bonds + present value of interest payments

Present value of bonds = face value x (P/F, 6%, 10)

N = 5 x 2 semi-annual periods = 10

Effective interest rate = 12% x 6/12 = 6%

Face value of bonds = $500,000

Present value of bonds = 500,000 x (P/F, 6%, 10)

= 500,000 x 0.5584 = $279,200

Present value of interest payments –

Semi-annual interest payments = 500,000 x 10% x 6/12 = $25,000

Present value of interest payments = 25,000 x (P/A, 6%, 10)

= 25,000 x 7.36 = $184,000

Selling price of bonds = 279,200 + 184,000 = $463,200

Selling price of the bonds = $463,200

Discount on bonds payable = 500,000 – 463,200 = $36,800

  1. Amortization schedule (effective interest method)

Amortization Schedule

Date

Interest Payment

Interest Expense

Discount Amortization

Carrying value of bonds

Jan 1, 2020

-

-

-

4,63,200

June 30, 2020

25,000

27,792

$2,792

$465,992

Dec 31, 2020

$25,000

$27,960

$2,960

$468,952

June 30, 2021

$25,000

$28,137

$3,137

$472,089

Dec 31, 21

$25,000

$28,325

$3,325

$475,414

June 30, 22

$25,000

$28,525

$3,525

$478,939

Dec 31, 22

$25,000

$28,736

$3,736

$482,675

June 30, 23

$25,000

$28,961

$3,961

$486,636

Dec 31, 23

$25,000

$29,198

$4,198

$490,834

June 30, 24

$25,000

$29,450

$4,450

$495,284

Dec 31, 24

$25,000

$29,717

$4,717

$500,000

total

$250,000

2,58,276

$36,800

Note: interest expense = 6% x carrying value of bond

Discount amortization = interest expense – interest payment

Carrying value beginning carrying value + discount amortization

  1. Journal entries for 2020:

Date

Account Titles and Explanation

Debit

Credit

Jan 1, 2020

Cash

$463,200

Discount on Bonds Payable

$36,800

Bonds Payable

$500,000

(To record issue of bonds)

June 30, 2020

Interest Expense

$27,792

Discount on Bonds Payable

$2,792

Cash

$25,000

(To record semi-annual interest payment)

Dec 31, 2020

Interest Expense

$27,960

Discount on Bonds Payable

$2,960

Cash

$25,000

(To record second interest payment)

  1. Entry to record retirement of bonds on July 1, 2023 at $104:

Date

Account Titles and Explanation

Debit

Credit

July 1, 2023

Bonds Payable

$500,000

Loss on Retirement of Bonds

$33,365

Cash

$520,000

Discount on Bonds Payable

$13,365

Computations:

Book value of bond at July 1, 2023 = face value – unamortized discount

Face value = 500,000

Unamortized discount on bonds payable = 13,365 (3,961 + 4,198 + 4,450 + 4,717)

Bonds net carrying value = $486,635

Reacquired value = $500,000 x 104% = 520,000

Loss on bond reacquisition = 520,000 – 486,635 = $33,365

  1. Straight line method of discount amortization –

Entry on June 30, 2020:

Date

Account Titles and Explanation

Debit

Credit

June 30, 2020

Interest Expense

$28,680

Discount on Bonds Payable

$3,680

Cash

$25,000

(To record semi-annual interest payment)

Computations:

Straight line method of discount amortization –

Discount on bonds payable = $36,800

Period = 10

Discount amortization for each semi-annual period = 36,800/10 = $3,680

Payment of interest = $25,000

Interest expense = 28,680